# Diagnosing the AGN population origin of TeV neutrinos with their spatial correlation

### v1

@@ VI.2 X-ray flux-weighting
Our main assumption is a linear relationship between the number of neutrinos and the AGN X-ray flux. Even if this linearity holds, different AGN classes may exhibit distinct proportionality constants due to underlying physical differences (Jacobsen2015; Berezinsky2024). To generalize this idea, we explore a family of models in which the neutrino yield from each AGN scales with its X-ray flux as $N_{\nu}\propto F_{X}^{s}$. The parameter $s$ reflects the assumed physical correlation strength: $s=0$ corresponds to no dependence (i.e., uniform weighting across AGNs), while increasing $s$ emphasizes the role of brighter AGNs in neutrino production. In this case, the source weights in Equation 2 of the likelihood function are modified to $w_{\alpha}={F_{\alpha}^{s}}/{\sum_{k}F_{k}^{s}}$ ensuring that the simulated neutrino–AGN associations reflect the chosen correlation index $s$.
Figure 6 (right) demonstrates the signal-to-noise ratio $(S/N)_{f}$ as a function of $s$ using a 30 km3 neutrino telescope over three years of exposure. The results indicate a monotonic increase in $(S/N)_{f}$ with $s$, reflecting the growing dominance of bright AGNs in neutrino production. For $s\approx 0$, corresponding to uniform weighting, the detection significance is limited to $\sim 3\sigma$. As $s$ increases toward unity—implying a direct proportionality between neutrino and X-ray flux—the significance improves substantially. This trend suggests that if neutrino emission preferentially traces the most luminous AGNs, the detectability of the population-level correlation is considerably enhanced. Conversely, a flatter weighting ($s<0.5$) implies that fainter AGNs contribute more evenly, reducing the overall signal contrast relative to the isotropic background.
VI.3 Blazars Neutrino Predictions
@@
Figure 7: Left: Cumulative fraction of the total neutrinos contributed by blazars with flux above a given threshold. The solid line shows the median, while the shaded regions indicate the 68% and 95 % credible intervals. Right: $(S/N)_{f}$ for the cumulative blazars contribution above 100 TeV, evaluated as a function of the assumed blazars fraction of the diffuse neutrino flux with a 30 km3 undersea neutrino telescope with 3 years of exposure.
@@
Blazars are also thought to be possible sources of extragalactic neutrinos, and are likely to contribute a significant fraction in the diffuse neutrino intensity at sub-PeV energies (e.g., Murase2014blazar; 2018SciTXS0506; Buson2022blazar). Motivated by the association of TXS 0506+056 with a hard-spectrum ($\Gamma\approx 2$) neutrino flare, we focus on the contribution of blazars to neutrinos above $\sim 100$ TeV, where such sources are expected to be more prominent.
@@
To directly test this scenario, IceCube has performed dedicated searches for cumulative emission from GeV-selected blazar populations (ICblazar2017ApJ). No significant excess was observed, and the upper limits constrain the contribution of 2LAC blazars to $\lesssim 27\%$ of the observed diffuse neutrino flux between 10 TeV and 2 PeV (for a spectral index $\Gamma=2.5$). Nevertheless, it is reasonable to expect that their relative contribution becomes more significant above $\sim 100$ TeV, motivating us to focus on this higher-energy regime in assessing the role of blazars as neutrino sources.
@@
We consider the 3933 classified as blazars from the fourth Fermi-LAT catalog (4FGLDR4), which includes 820 flat spectrum radio quasars (FSRQs), 1490 BL Lacs, and 1623 other sources classified as ”blazars of unknown type”. We extrapolate their collective neutrino contribution using the GeV gamma-ray luminosity function (GLF) (BlazarsGLF2020PhRvD), following the same procedure as that applied to the AGN X-ray luminosity function (XLF) in Section V. Figure 7 (left) presents the cumulative fraction of the total blazar neutrino flux as a function of flux threshold, with shaded regions showing the 68% and 95% credible intervals. Here, the “flux threshold” corresponds to the average energy flux detection limit of Fermi-LAT, representing the approximate completeness threshold of the blazar sample. Given that the mean sensitivity limit for the Fermi-LAT blazar sample is about $2\times 10^{-12}\,\mathrm{erg\,s^{-1}\,cm^{-2}}$ (fermicat2022ApJS), we estimate that this Fermi-LAT blazar population accounts for approximately $\sim 85\%\pm 5\%$ of the total neutrino emission from all blazars. Figure 7 (right) illustrates the trend of the detection significance as a function of the assumed blazars fraction for a 30 km3 detector over 3 years.
@@
When systematic effects are included, the dominant limitation arises from GLF uncertainties. In this case, if blazars contribute to $\sim$5% of the diffuse neutrino flux above 100 TeV, our method reaches a $5\sigma$ detection significance for 3 years of exposure. For larger assumed fractions, the significance does not further increase, since the GLF uncertainties impose a systematic floor on the achievable sensitivity. This demonstrates that our approach is capable of identifying even a modest blazar contribution to the diffuse neutrino flux with high confidence. In particular, the ability to reach a $5\sigma$ significance level for a $\sim$5% contribution highlights the strong discovery potential of the method.
VI.4 Implications of a Two-Population AGN Scenario
The current evidence suggests that different classes of AGNs may contribute neutrinos in distinct energy bands. The case of NGC 1068 (NGC1068_ic_2022Sci) indicates that at least some Seyfert galaxies predominantly produce neutrinos at relatively low energies ($E_{\nu}<100$ TeV). In contrast, the presence of >100>100 TeV neutrinos in the direction of NGC 7469 (although not yet statistically significant, NGC7469-2025ApJ-2nu) or TXS 0506+056 (2018SciTXS0506) hints at the possibility that another subclass of AGNs could be responsible for the high-energy tail of the diffuse neutrino spectrum (Yang et al. in preparation). These considerations motivate a two-population AGN scenario, in which (i) NGC 1068-like sources dominate the low-energy neutrino flux, while (ii) a different population, potentially blazar-like or luminous Seyferts with favorable conditions for hard neutrino spectrum, contributes primarily to the >100>100 TeV neutrino flux. The diffuse astrophysical neutrino background would then emerge as a superposition of these two populations.
We further examined the simplified hypothesis that neutrinos with $E_{\nu}>100$ TeV are predominantly produced by a subset of the AGN population (’hard’ AGN). Since the physical origin of such hard-spectrum AGNs is not yet identified, we randomly select a given fraction (e.g., 25% or 50%) of the AGN sample to represent this subpopulation in our simulations. Compared to the assumption that the entire AGN population contributes to the diffuse flux, restricting the contribution to only a fraction of AGNs effectively increases the weight assigned to each source in the likelihood analysis. Such hard-spectrum AGNs may feature higher proton acceleration efficiencies or denser photon fields, enhancing $p\gamma$ interactions at higher energies. This reduced source population strengthens the contrast between signal and background, thereby enhancing the achievable confidence level of detection. In the future, more refined AGN neutrino models may allow us to classify sources based on physical parameters that naturally lead to either soft or hard neutrino spectra. Such developments would provide a critical pathway to test whether a small subset of hard AGNs indeed dominates the diffuse neutrino flux above 100 TeV.
At the same time, it should be noted that the dominant source of systematic uncertainty remains in the X-ray luminosity function of AGNs, as discussed in Sec. V. This implies that the ultimate confidence level for testing the hard-AGN hypothesis is limited by the same luminosity-function uncertainties as in the baseline analysis. Nevertheless, the method remains robust and provides a promising avenue to probe whether a fraction of AGNs with harder spectra are the main contributors to the diffuse neutrino flux above 100 TeV.
VII Conclusion
@@
We have presented a comprehensive framework to evaluate the detectability of correlations between high-energy neutrinos and AGNs, incorporating both statistical and systematic effects. Our analysis combines flux-weighted likelihood methods with realistic detector performance and observational constraints. The key findings are summarized as follows:
@@
  • Dominant role of angular resolution: The sensitivity of AGN–neutrino correlation searches is primarily driven by the detector’s point-spread function. Achieving sub-degree resolution, particularly $\lesssim 0.1^{\circ}$, dramatically reduces background contamination and enhances detection significance. For example, with $2000$ events and a $0.1^{\circ}$ resolution, a $5\sigma$ detection is achievable, whereas resolutions coarser than $0.4^{\circ}$ would require orders of magnitude more events.
@@
  • Source separation criterion: To minimize source confusion and ensure robust association between neutrinos and AGNs, we require the mean angular separation between selected AGNs to exceed three times the detector’s PSF width ($\bar{d}>3\sigma_{\rm psf}$). This criterion effectively avoids source confusion, yielding a cleaner AGN sample for correlation analysis and more reliable flux extrapolations to the full population.
@@
  • Energy threshold optimization: Increasing the minimum neutrino energy threshold improves angular resolution and suppresses atmospheric backgrounds, but reduces event statistics. For a 30 km3 undersea neutrino telescope (e.g., HUNT) that features rapidly improving resolution with energy (down to $\sim 0.04^{\circ}$ above 300 TeV), the optimal sensitivity is reached at thresholds of a few hundred TeV. In contrast, IceCube-Gen2 shows weaker energy dependence in its resolution; for this detector, selecting $E_{\nu}>100$ TeV strikes a better balance between angular precision and event statistics.
  • Flux-limited AGN extrapolation and XLF systematics: Our likelihood analysis is performed on a flux-limited AGN subset ($\bar{d}>3\sigma_{\rm psf}$), and the total neutrino contribution is inferred via extrapolation to the full AGN population. The extrapolation uncertainty is dominated by the flux threshold $F_{\rm min}$: for $F_{\rm min}\lesssim 10^{-14}\ {\rm erg\ cm^{-2}\ s^{-1}}$ the uncertainty is small ($\lesssim 10\%$), while $F_{\rm min}\gtrsim 3\times 10^{-13}\ {\rm erg\ cm^{-2}\ s^{-1}}$ yields unreliable results due to the dominance of faint AGNs in the XLF. In addition, accounting for realistic distributions of reconstructed event angles introduces an additional $\sim 10\%$ uncertainty in the inferred AGN neutrino fraction $f$, particularly affecting low-energy dominated samples. This highlights the importance of incorporating per-event angular information in future analyses.
  • Implications for next-generation neutrino telescopes: In the above 1300 TeV energy range, a 30 km3 undersea neutrino telescope is expected to reach a significance of $8.6\sigma$ after five years of exposure. The limited improvement despite the larger statistics reflects the dominant impact of XLF extrapolation uncertainties, which ultimately set a systematic floor on the achievable sensitivity. For IceCube-Gen2, with $\sim 0.2^{\circ}$ resolution above 100 TeV, the achievable confidence level is $\sim 2.4\sigma$ over five years.
@@
These results demonstrate that next-generation neutrino observatories with sub-degree angular precision and sufficient exposure have the potential to either confirm or strongly constrain the AGN origin of high-energy neutrinos. Deep multi-wavelength AGN surveys, especially in X-rays, are crucial to minimize uncertainties in flux extrapolation and to fully exploit the discovery potential of upcoming facilities.
We thank Tian-Qi Huang and Siqi Yu for valuable discussions.

### v3

@@ V.2 X-ray flux-weighting
So far, we have assumed a linear relationship between the number of neutrinos and the AGN X-ray flux. This treatment considers that the X-ray flux of an AGN may be a proxy for its high-energy proton power and also representative for the target density for neutrino production. However, the linear relationship is unwarranted yet (e.g., Ambrosone, 2024; Saurenhaus et al., 2025). To generalize this idea, we explore a family of models in which the neutrino flux from an AGN scales with its X-ray flux as $N_{\nu}\propto F_{X}^{s}$. The parameter $s$ reflects the assumed physical correlation strength: $s=0$ corresponds to no dependence (i.e., uniform weighting across AGNs), while increasing $s$ emphasizes the role of brighter AGNs in neutrino production. In this case, the source weights in Equation 5 of the likelihood function are modified to $w_{j}={F_{j}^{s}}/{\sum_{k}F_{k}^{s}}$ ensuring that the simulated neutrino–AGN associations reflect the chosen correlation index $s$.
We also consider a special scenario in which all astrophysical neutrinos originate from AGNs, corresponding to $\alpha=1$. Figure 8 summarizes the impact of the flux-weighting parameter $s$ on both the precision of the recovered AGN neutrino fraction and the overall detection significance, for a 30 km3 neutrino telescope with three years of exposure. The top panel of Figure 8 displays the relative uncertainty of the recovered AGN neutrino fraction $\delta$. We find that $\delta$ decreases monotonically with increasing $s$, indicating that the precision of the recovered AGN contribution improves as neutrino emission becomes more strongly associated with X-ray–bright sources. The bottom panel shows the detection significance, quantified by $\sqrt{\rm TS}$, as a function of $s$. A clear monotonic increase is observed, indicating that stronger weighting toward luminous AGNs substantially enhances the detectability of the population-level correlation. In the limit $s\simeq 0$, corresponding to uniform source weighting, the detection significance remains modest, at the level of $\sim 2\sigma$. As $s$ approaches unity, implying a near-linear proportionality between neutrino and X-ray flux, the significance improves markedly, reflecting the growing dominance of the brightest AGNs in the signal. This behavior can be understood as a consequence of two related effects. First, larger values of $s$ effectively suppress the contribution from faint AGNs while simultaneously increasing the relative weight of X-ray–bright sources, which significantly enhances the detection significance and reduces the statistical uncertainty in the recovered AGN neutrino fraction. Second, smaller values of $s$ assign a relatively larger weight to faint AGNs, for which the extrapolation of the XLF introduces substantial systematic uncertainties, further inflating the uncertainty in the recovered fraction.
Overall, these results suggest that if astrophysical neutrino production preferentially traces the most luminous AGNs, both the detectability of the AGN–neutrino correlation and the precision of the inferred AGN neutrino fraction are significantly improved. Conversely, weaker flux weighting ($s<0.5$) implies a larger relative contribution from faint AGNs, leading to reduced signal contrast with respect to the isotropic background and increased uncertainty in the recovered fraction.
V.3 Blazars Neutrino Predictions
@@
Figure 9: Left: Cumulative fraction of the total neutrinos contributed by blazars as a function of the gamma-ray flux threshold of blazars. The solid line shows the median, while the shaded regions indicate the 68% and 95 % credible intervals. Right: Top panel: Same as the Figure 7, recovered blazar neutrinos fraction $f_{\rm tot}$ as a function of the blazar neutrino fraction $\alpha_{\rm blazars}$. Bottom panel: Corresponding detection significance $\sqrt{\rm TS}$ as a function of $\alpha_{\rm blazars}$, showing that even partial blazars contributions can be detected with high statistical significance.
@@
Blazars are also widely discussed as possible sources of extragalactic neutrinos, and are likely to contribute a significant fraction in the diffuse neutrino intensity at sub-PeV energies (e.g., Murase et al. 2014; IceCube Collaboration et al. 2018; Das et al. 2022; Buson et al. 2022). Motivated by the association of TXS 0506+056 with a hard-spectrum ($\Gamma\approx 2$) neutrino flare, we focus on the contribution of blazars to neutrinos above $\sim 100$ TeV, where such sources are expected to be more prominent.
@@
To directly test this scenario, IceCube has performed dedicated searches for cumulative emission from GeV-selected blazar populations (Aartsen et al., 2017). No significant excess was observed, and the upper limits constrain the contribution of 2LAC blazars to $\lesssim 27\%$ of the observed diffuse neutrino flux between 10 TeV and 2 PeV (for a spectral index $\Gamma=2.5$). Nevertheless, it is reasonable to expect that their relative contribution becomes more significant above $\sim 100$ TeV, and hence we focus on this higher-energy regime to assess the role of blazars as neutrino sources.
@@
We consider 3933 blazars from the fourth Fermi-LAT catalog (Ballet et al., 2023), which provide a representative sample for our study. We also extrapolate their collective neutrino contribution using the GeV gamma-ray luminosity function (GLF) (Manconi et al., 2020), following the same procedure as that applied to the AGN XLF in Section IV. The left panel of Figure 9 presents the cumulative fraction of the total blazar neutrino flux as a function of flux threshold, with shaded regions showing the 68% and 95% credible intervals. Here, the “flux threshold” corresponds to the average energy flux detection limit of Fermi-LAT, representing the approximate completeness threshold of the blazar sample. Given that the mean sensitivity limit for the Fermi-LAT blazar sample is about $2\times 10^{-12}\,\mathrm{erg\,s^{-1}\,cm^{-2}}$ (Ajello et al., 2022), we estimate that this Fermi-LAT blazar population accounts for approximately $\sim 85\%\pm 5\%$ of the total neutrino emission from all blazars. This result is broadly consistent with previous Fermi-LAT studies of the blazar population at high energies. In particular, analyses of the extragalactic $\gamma$-ray background above $\sim 50$ GeV have shown that the bulk of the total $\gamma$-ray energy density is dominated by resolved blazars, while the contribution from unresolved, low-flux sources is subdominant (e.g., Ackermann et al., 2016). Similar conclusions have been reached through stacking analyses of faint and extreme blazar populations, which indicate that unresolved sources contribute only a modest fraction of the total $\gamma$-ray flux (Paliya et al., 2019).
@@
The right panel of Figure 9 illustrates the trend of the detection significance as a function of the assumed blazars fraction for the HUNT, a 30 km3 detector over 3 years, with systematic uncertainties included. When systematic effects are included, the dominant limitation arises from GLF uncertainties. In this case, if blazars contribute to $\sim$5% of the diffuse neutrino flux above 300 TeV, our method reaches a $\sqrt{\rm TS}>9$ detection significance for 3 years of exposure. For larger assumed fractions, the significance saturates, since the GLF uncertainties impose a systematic floor on the achievable sensitivity. This demonstrates that our approach is capable of identifying even a modest blazar contribution to the diffuse neutrino flux with high confidence.
VI Conclusions
@@
We have evaluated the prospect of diagnosing the correlation between X-ray AGNs and high-energy neutrinos with future neutrino telescopes, incorporating both statistical and systematic effects. By combining a flux-weighted likelihood method with realistic detector performance and observational constraints, we have found that:
@@
  • Dominant role of angular resolution: The sensitivity of AGN–neutrino correlation is primarily driven by the point-spread function of neutrino detectors. Achieving sub-degree resolution, particularly $\lesssim 0.1^{\circ}$, dramatically reduces background contamination and enhances the significance of detection. For example, with $2000$ events and a $0.1^{\circ}$ resolution, a $5\sigma$ detection is achievable, whereas resolutions coarser than $0.4^{\circ}$ would require orders of magnitude more events.
@@
  • Source separation criterion: To minimize source confusion and ensure robust association between neutrinos and AGNs, we require the mean angular separation between two selected AGNs to exceed three times of the detector’s PSF width (i.e., $\bar{d}>3\sigma_{\rm psf}$). This criterion effectively avoids source confusion.
@@
  • Energy threshold optimization: Increasing the minimum neutrino energy threshold improves the angular resolution and suppresses the atmospheric neutrino background, but at the cost of reducing the event statistics. For a 30 km3 undersea/underwater neutrino telescope (e.g., HUNT) that features rapidly improving angular resolution with energy (down to $\sim 0.04^{\circ}$ above 300 TeV), the optimal sensitivity is reached at a threshold of a few hundred TeV. In contrast, IceCube-Gen2 shows a weaker energy dependence in its angular resolution; for this detector, selecting $E_{\nu}>100$ TeV strikes a better balance between the angular precision and event statistics.
  • Flux-limited AGN extrapolation and XLF systematics: Our likelihood analysis is performed on a flux-limited AGN subset selected to satisfy $\bar{d}>3\sigma_{\rm psf}$, ensuring that source confusion does not bias the inference. The total neutrino contribution from the full AGN population is then obtained via extrapolation using the AGN XLF. The associated uncertainty is strongly controlled by the adopted flux threshold $F_{\rm min}$. For $F_{\rm min}\leq 10^{-14}\,{\rm erg\ cm^{-2}\ s^{-1}}$, the extrapolation uncertainty remains small ($\leq 10\%$), whereas for $F_{\rm min}\geq 7\times 10^{-13}\,{\rm erg\ cm^{-2}\ s^{-1}}$ the inferred total contribution becomes unreliable, as the correction factor is dominated by uncertainties in the faint-end slope and normalization of the XLF.
  • Implications for next-generation neutrino telescopes: In the energy range above 300 TeV, a 30 km3-scale underwater neutrino telescope is expected to reach a significance of $8.7\sigma$ with one year of exposure. With five years of exposure, the significance increases to $\sim 19\sigma$, while the total AGN neutrino fraction can be constrained to a relative precision of $\delta\simeq 10\%$. For IceCube-Gen2, considering all neutrinos above 100 TeV, the achievable confidence level is $\sim 3.1\sigma$ with five years of exposure.
  • Sensitivity to partial AGN contributions: For a 30 km3-scale underwater neutrino telescope, if AGNs contribute a fraction $\alpha\leq 0.17$ of the total diffuse neutrino flux, it is not distinguishable from an isotropic background at the $3\sigma$ level with three years of exposure. Even with an extended exposure of 10 years, contribution below $\alpha\simeq 0.09$ remains undetectable at the same significance. For a larger contribution, the detection significance scales approximately linearly with $f_{\rm tot}$.
@@
These results demonstrate that next-generation neutrino observatories with sub-degree angular precision and sufficient exposure have the potential to confirm or strongly constrain the AGN origin of high-energy neutrinos. Deep multi-wavelength AGN surveys, especially in X-rays, are crucial to minimize uncertainties in the flux extrapolation and to fully exploit the discovery potential of upcoming facilities.
We would like to thank Tian-Qi Huang and Shiqi Yu for valuable discussions. This work is supported by the National Natural Science Foundation of China (grant Nos. 12333006, 12121003 and 12393852), National SKA Program of China (2025SKA0110104) and the Fundamental Research Funds for the Central Universities (KG202502). We are grateful to the High Performance Computing Center of Nanjing University for doing the numerical calculations in this paper on its blade cluster system.
Appendix A Background TS Distribution and Significance Calibration
[…]
To validate the use of $\sqrt{\mathrm{TS}}$ as an approximate significance estimator,we generate $5\times 10^{6}$ background-only Monte Carlo realizations for the HUNT detector configuration.For each realization, the test statistic (TS) is computed in the same way as for the observed data.
[…]
As seen in the figure, the background TS distribution is in good agreement with the Wilks expectation for the TS range of interest,justifying the common approximation of statistical significance as $\sigma\simeq\sqrt{\mathrm{TS}}$. Small deviations may occur at very low TS values, but they do not affect the interpretation of the results in the parameter space considered.